OpenStudy (bloomlocke367):

Graph the conic section. 25x^2-16y^2=400. I'll fan and medal. It's a hyperbola, I know that, but I don't know what to do with it..

OpenStudy (bloomlocke367):

@ganeshie8

OpenStudy (anonymous):

use desmos.com it graphs it all for you and tells you the points :)

OpenStudy (bloomlocke367):

but I still would like an explanation on how to do it..

OpenStudy (anonymous):

okay hold on. let me work this out.

OpenStudy (anonymous):

are you wanting to solve the equation too?

OpenStudy (bloomlocke367):

yes..

OpenStudy (anonymous):

okay. well to solve the equation. you're going to divide 400 by both sides. 25x^2/400-16y^2/400=1 …. this is in standard form of a hyperbola [x^2/a^2-y^2/b^2=1] which intersects the x-axis at (20/5,0) and (-20/5,0). there are two asymptotes: y=(400/16)/(400/25)*x and y = - (400/16)/(400/25)*x

OpenStudy (anonymous):

do you understand that.. or need further elaboration?

OpenStudy (bloomlocke367):

further explanation... please. what's asymptotes?

OpenStudy (anonymous):

thats a line that continually approaches a curve but doesn't meet at any finite distance

OpenStudy (anonymous):

all the work i gave you is everything you need to graph it.. would u like me to show you the graph and what it looks like?

OpenStudy (bloomlocke367):

no, I got it. thanks :) how did you find the intercepts?

OpenStudy (anonymous):

no problem:) intercepts of what?

OpenStudy (anonymous):

hyperbola?

OpenStudy (bloomlocke367):

you said it intercepts the x-axis at (4,0) and (-4,0).. how'd you get that?